# Multiple dispersive bounds. I) An improved z-expansion

### v2

@@ Abstract
We propose the implementation of two novel ingredients in the standard $z$-expansion of hadronic form factors, commonly referred to as the Boyd-Grinstein-Lebed (BGL) approach [1, 2, 3, 4]. The first new ingredient is the explicit addition of a unitarity filter applied to a given known set of input data for the hadronic form factors. This further constraint is not usually taken into account in the phenomenological applications of the BGL $z$-expansion. We show that it brings to a formulation of the BGL approach fully equivalent to the Dispersion Matrix (DM) method [5], which describes hadronic form factors in a completely model-independent and non-perturbative way. The second key ingredient is represented by the introduction of suitable kernel functions in the evaluation of unitarity bounds, leading to the application of multiple dispersive bounds to hadronic form factors, whenever data and/or (non-)perturbative techniques allow to do so. This idea may be useful for the investigation of many physical processes, from the analysis of the electromagnetic form factors of mesons and baryons to the study of weak semileptonic decays of hadrons. An explicit numerical application will be presented in the companion paper [6], where the effects of sub-threshold branch-cuts are analyzed.
@@ I Introduction
In the present work and in the companion one [6], we propose two main improvements of the BGL expansion usually adopted in the literature.
@@
This first paper is organized as follows. In Section II we review the basic properties of the BGL $z$-expansion and, then, discuss the different constraints dictated by unitarity both on the coefficients of the BGL expansion and on the given input dataset for the form factors. To this end we derive an additive representation of the BGL expansion, which turns out to be equivalent to the prediction of the DM approach. In Section III, we introduce the idea of multiple dispersive bounds and discuss how they can be implemented in the BGL expansion in absence of sub-threshold branch-cuts. The case in which sub-threshold branch-cuts are present is analyzed separately in the companion paper [6]. Our conclusions are summarized in Section IV, while Appendices A and B contain some details concerning aspects of the unitary BGL framework.
@@ II The BGL expansion
In this Section we briefly recall some basic features of the BGL $z$-expansion. For sake of simplicity, we consider the case of a single, generic form factor $f(t)$, where $t$ is the four-momentum transfer, describing the transition induced by a current $J$ between two hadrons $H_{1}$ and $H_{2}$ with masses $m_{1}$ and $m_{2}$, respectively.
@@
The BGL expansion of the form factor $f(z)$ onto the unit disk reads as [4]
@@ II.1 The dispersive susceptibility as the upper bound $\chi^{U}$
To be more specific, the idea is to consider the polarization function $\Pi(q^{2})$ associated to a double insertion of the current $J$. The polarization function satisfies a dispersion relation and its imaginary part on the branch-cut, dictated by unitarity, is given by a sum over a complete set of on-shell states $\overline{H}_{1}H_{2}$ that couple the current $J$ to the vacuum. Using crossing symmetry and the positivity of each term in the completeness sum one arrives to a bound for the form factor $f(t)$ of the form [4]
@@
where $W_{n}(t)$ is a known weight function depending on the particular form factor $f(t)$ and $\chi_{n}$ is a derivative of the polarization function, the so-called susceptibility, namely (cf., e.g., Ref. [5])
@@
In the case of the form factors relevant for the semileptonic decays $M_{1}\to M_{2}\ell\nu_{\ell}$, with $M_{1}$ being a pseudoscalar meson and $M_{2}$ a pseudoscalar or vector meson, the final expressions for the kinematical function $\phi$ at $q_{0}^{2}=0$ reads as [4, 29]
@@
where $n_{I}$ is an isospin Clebsch-Gordan factor, the constant $\delta$ depends on the precise definition of the form factor and the values of the parameters $K$, $a$, $b$, $c$ and $n$ are summarised in Table 1 of Appendix A in the case of the the form factors relevant for the semileptonic decays $M_{1}\to M_{2}\ell\nu_{\ell}$. When the susceptibility $\chi_{n}(q_{0}^{2})$ is evaluated at a non-vanishing value of $q_{0}^{2}$, the function $\phi$ must be multiplied by an additional factor [4], which, thanks to Eq. (10), is given by $t/(t-q_{0}^{2})]^{(n+1)/2}$ on the unit circle. According to Eqs. (11) its extension inside the unit disk leads to the final factor
@@
The basic step to develop such non-perturbative computations is to consider the Euclidean current-current correlator $V(\tau)$ depending on the Euclidean time-distance $\tau$ (i.e. the insertion of two currents $J$ integrated over their spatial separation). Such a correlator is related to the polarization function $\Pi(Q^{2})$ evaluated at an Euclidean four-momentum $Q^{2}\equiv-q^{2}$ by
@@ II.2 Unitary BGL expansion
One of the main goals of this work is to show that, in reality, in order to be consistent with unitarity there is a further hidden constraint beyond the usual one given by Eq. (7), or by Eq. (21) in the truncated version, that should be taken into account. To this end, let us consider the set of functions represented by all the BGL expansions (4) consistent both with the unitarity constraint (7) and with a given set of input data $\{f_{i}\}$. For sake of simplicity, in what follows we assume that the the form factor $f(z)$ is known for a series of distinct real values of $t=t_{i}<t_{+}$ (with i=1,2,,Ni=1,2,...,N), so that all the $z_{i}$ are real and different from each other. The untruncated BGL expansion (4) of the form factor $f(z)$ can be rewritten exactly in an additive form. Here below we use directly the final result of the computation carried out in Appendix B, namely
@@
The interpretation of the inequalities given by Eqs. (25-26) is as follows. The new inequality (25) is necessary to guarantee that the function (23) is consistent with unitarity and the remainder function (24) is analytic of the real type. Eq. (25) represents a filter acting directly on the input data and allows to select only the subset $\{\overline{f}_{i}\}$ of input data, which can be reproduced by BGL expansions satisfying the unitarity constraint (7). Furthermore, the inequality (26), acting on the coefficients $c_{k}$ of the remainder function $R(z)$, is a consequence of both the unitarity constraint $\chi[f]\leq\chi^{U}$, which is equivalent to the inequality (7) on the coefficients $a_{k}$ of the BGL expansion, and the new inequality (25). Indeed, an explicit calculation, shown in Appendix B, yields
@@
We stress that the above result implies two different unitarity constraints. The first one is given by $\chi[f]\leq\chi^{U}$, which implies the unitarity constraint (7), which guarantees that all the BGL expansion (22) are unitary. The other constraint is given by $\chi[\beta]\leq\chi^{U}$, which guarantees that the function (23) fulfills unitarity. It tells us whether a given set of input data $\{f_{i}\}$ is consistent with unitarity, i.e. whether it may be reproduced exactly by a BGL expansion (4) with coefficients $a_{k}$ satisfying the unitarity constraint (7). When Eq. (25) is not fulfilled, the input data $\{f_{i}\}$ are not consistent with unitarity and should be removed from any analysis since they contain non-unitary effects [^4] A consequence of the presence of non-unitary input data in a fitting procedure based on the BGL $z$-expansion is the value of the $\chi^{2}$-variable. Indeed, the latter one may receive both a first contribution related to the truncation order $M$ in Eq. (20) and a second one due to the fact that non-unitary input data cannot be reproduced by a unitary BGL expansion. The latter contribution cannot be eliminated by increasing the order $M$ of the truncation and it is typically associated to a saturation of the dispersive bound $\chi^{U}$, i.e. $\chi[f]\simeq\chi^{U}$.. This may happen in various circumstances: when uncorrelated datasets of input data should be combined and/or when the correlations among input data are not determined precisely taking properly into account unitarity, as e.g. in the case of the semileptonic $B\to D^{*}$ form factors discussed extensively in Ref. [36]. In all cases the application of the filter (25) allows to reconstruct the proper covariance matrix of the input data consistent with unitarity. In this respect, we mention the importance sampling procedure, developed in Ref. [34], which must be adopted in the case of a large number of input data points for a proper application of the unitarity filter $\chi[\beta]\leq\chi^{U}$.
Thus, in order to summarize the results of this Section a fitting procedure based on a unitary BGL $z$-expansion of the form factor $f(z)$ should be made of the following steps.
@@
    Check the sample of input data $\{f_{i}\}$ generated according to a given original covariance matrix $C_{ij}$ against the unitarity filter $\chi[\beta]\leq\chi^{U}$.
@@
    Select only the subset of input data $\{\overline{f}_{i}\}$ with a corrected covariance matrix $\overline{C}_{ij}$ satisfying the unitarity filter $\chi[\beta]\leq\chi^{U}$.
@@
    Apply the truncated BGL expansion (20) to the unitary subset $\{\overline{f}_{i}\}$ and impose on the coefficients $a_{k}^{(M)}$ the unitarity constraint (21).
@@
which is nothing else than the final result of the DM method of Ref. [5], originally proposed in Refs. [27, 37]. Without relying on any truncation or fitting procedure, Eq. (33) represents the uniform distribution of values of the form factor $f(z)$ at a generic value of $z$ consistent with unitarity as well as with the input data $\{\overline{f}_{i}\}$. We stress that in the DM approach the unitarity filter $\chi[\beta]\leq\chi^{U}$ on the input data is always taken properly into account, since $\gamma(z)$ must be non-negative. The impact of the unitarity filter $\chi[\beta]\leq\chi^{U}$ has been addressed in the case of the semileptonic $B_{(s)}\to D_{(s)}^{(*)}\ell\nu_{\ell}$ decays in Refs.[5, 38, 36, 39, 14], for the semileptonic $B_{(s)}\to\pi(K)\ell\nu_{\ell}$ decays in Ref. [33] and for the pion electromagnetic form factor in Ref. [34]. We highlight that the inclusion of the unitarity filter $\chi[\beta]\leq\chi^{U}$ may lead to important reduction of the uncertainties associated to the extrapolation of the form factors, especially in presence of a large number of input data (see, e.g., Fig. 6 of Ref. [14] in the case of the semileptonifc $B\to D^{*}\ell\nu_{\ell}$ form factors). We stress that an interesting generalization of Eq. (29), which includes also the first $K$ derivatives of the form factor at the origin $z=0$, was derived in Ref. [40] (see also Ref. [41]).
Finally, we comment briefly on the role of the auxiliary variable $t_{0}$, which fixes the value of $t$ at which $z(t_{0};t_{0})=0$. The form factor $f$ cannot depend on the choice of $t_{0}$. Thus, since the Blaschke product $B(z)$ as well as the bound $\chi^{U}$ are independent of $t_{0}$ by construction, both the kinematical function $\phi(z)$ and the BGL coefficients $a_{k}$ must depend on $t_{0}$ in such a way to cancel any possible dependence of $f(z)$ on $t_{0}$. This implies that any difference in the truncated BGL $z$-expansions (20) corresponding to different values of $t_{0}$ may signal that the truncation error is not under full control. Alternatively, the value of $t_{0}$ can be chosen according to the requirement of accelerating as much as possible the convergence of the truncated BGL expansion (20). In Refs. [1, 3, 42] it is argued that for semileptonic process $H_{1}\to H_{2}\ell\nu_{\ell}$ an optimal choice could be $t_{0}=t_{opt}=(m_{1}+m_{2})\left(\sqrt{m_{1}}-\sqrt{m_{2}}\right)^{2}$, which minimizes the truncation error of the expansion (20) estimated as in Refs. [3, 4].

### v3

# Multiple dispersive bounds. I) The z-expansion

@@ Abstract
We propose the implementation of two ingredients in the phenomenological applications of the unitary approach based on the $z$-expansion of hadronic form factors, commonly referred to as the Boyd-Grinstein-Lebed (BGL) $z$-expansion [1, 2, 3, 4]. The first ingredient is the explicit addition of a unitarity filter applied to a given set of input data for the hadronic form factors. This further constraint is not usually taken into account in the phenomenological applications of the BGL $z$-expansion. We show that it follows from the equivalence between the BGL approach and the Dispersion Matrix (DM) method [5], which also describes hadronic form factors in a completely model-independent and non-perturbative way. The second ingredient is represented by the introduction of suitable kernel functions in the evaluation of unitarity bounds, leading to the application of multiple dispersive bounds to hadronic form factors, whenever data and/or (non-)perturbative techniques allow to do so. This idea may be useful for the investigation of many physical processes, from the analysis of the electromagnetic form factors of mesons and baryons to the study of weak semileptonic decays of hadrons. An explicit numerical application will be presented in the companion paper [6], where the effects of sub-threshold branch-cuts are analyzed.
@@ I Introduction
In the present work and in the companion one [6], we propose two main improvements of the unitary analysis, based on the BGL $z$-expansion, usually adopted in the literature for reproducing a given set of input data.
@@
This first paper is organized as follows. In Section II we review the basic properties of the BGL $z$-expansion and, then, discuss the different constraints dictated by unitarity both on the coefficients of the BGL expansion and on the given input dataset for the form factors. To this end we derive again an additive representation of the BGL expansion, which turns out to be equivalent to the prediction of the DM approach. In Section III, we introduce the idea of multiple dispersive bounds and discuss how they can be implemented in the BGL expansion in absence of sub-threshold branch-cuts. The case in which sub-threshold branch-cuts are present is analyzed separately in the companion paper [6]. Our conclusions are summarized in Section IV, while Appendices A and B contain some details concerning aspects of the unitary BGL framework.
@@ II The BGL expansion
In this Section we briefly recall some basic features of the BGL $z$-expansion. For sake of simplicity, we consider the case of a single, generic form factor $f(t)$, where $t$ is the squared four-momentum transfer, describing the transition induced by a current $J$ (which may be the weak VAV-A current or the electromagnetic vector current) between two hadrons $H_{1}$ and $H_{2}$ with masses $m_{1}$ and $m_{2}$, respectively.
@@
The BGL expansion of the form factor $f(z)$ on the unit disk reads as [4]
@@ II.1 The dispersive susceptibility as the upper bound $\chi^{U}$
To be more specific, the idea is to consider the polarization function $\Pi(q^{2})$ associated to a double insertion of the hadronic current $J$. For instance, in the case of the weak vector or axial flavor-changing currents the polarization tensor $\Pi_{\mu\nu}(q^{2})$ is defined as
$\Pi_{\mu\nu}(q^{2})$ $=$ $i\int d^{4}x\,e^{iq\cdot x}\,\langle 0|{\cal{T}}\{J_{\mu}^{\dagger}(x)J_{\nu}(0)\}|0\rangle$~ (9)
$=$ $q_{\mu}q_{\nu}\,\Pi_{L}(q^{2})+\left(q_{\mu}q_{\nu}-g_{\mu\nu}q^{2}\right)\,\Pi_{T}(q^{2})~,$~
where $\cal{T}$} is the time-ordered product and $\Pi_{L(T)}(q^{2})$ are the longitudinal(transverse) polarization functions. The polarization function $\Pi(q^{2})$, which may be $\Pi_{L}(q^{2})$ or $\Pi_{T}(q^{2})$ in Eq. (9), satisfies a dispersion relation and its imaginary part on the branch-cut, dictated by unitarity, is given by a sum over a complete set of on-shell states $\overline{H}_{1}H_{2}$ that couple the current $J$ to the vacuum. Using crossing symmetry and the positivity of each term in the completeness sum one arrives to a bound for the form factor $f(t)$ of the form [4]
@@
where $W_{n}(t)$ is a known weight function depending on the particular form factor $f(t)$. In Eq. (10) the quantity $\chi_{n}$ is a derivative of the polarization function, the so-called susceptibility, namely (cf., e.g., Ref. [5])
@@
In the case of the form factors relevant for the semileptonic decays $M_{1}\to M_{2}\ell\nu_{\ell}$, with $M_{1}$ being a pseudoscalar meson and $M_{2}$ a pseudoscalar or vector meson, the final expression for the kinematical function $\phi$ at $q_{0}^{2}=0$ reads as [4, 29]
@@
where $n_{I}$ is an isospin Clebsch-Gordan factor, the constant $\delta$ depends on the precise definition of the form factor and the values of the parameters $K$, $a$, $b$, $c$ and $n$ are summarised in Table 1 of Appendix A. When the susceptibility $\chi_{n}(q_{0}^{2})$ is evaluated at a non-vanishing value of $q_{0}^{2}$, the function $\phi$ must be multiplied by an additional factor [4], which, thanks to Eq. (11), is given by $t/(t-q_{0}^{2})]^{(n+1)/2}$ on the unit circle. According to Eqs. (12) its extension inside the unit disk leads to the final factor
@@
The basic step to develop such non-perturbative computations is to consider the Euclidean current-current correlator $V(\tau)$ depending on the Euclidean time-distance $\tau$ (i.e. the insertion of two currents $J$ integrated over their spatial separation)
$V(\tau)=\int d\vec{x}~\langle J^{\dagger}(\vec{x},\tau)J(\vec{0},0)\rangle~,$~ (16)
where $\langle$$\rangle$ means the average of the $T$-product of the two currents $J$ over gluon and fermionic fields. Such a correlator is related to the polarization function $\Pi(Q^{2})$ evaluated at an Euclidean squared four-momentum $Q^{2}\equiv-q^{2}$ by [34, 35]
@@ II.2 Unitary BGL expansion
One of the main goals of this work is to show that, in reality, in order to be consistent with unitarity there is a further hidden constraint beyond the usual one given by Eq. (7), or by Eq. (23) in the truncated version, that should be taken into account. To this end, let us consider the set of functions represented by all the BGL expansions (4) consistent both with the unitarity constraint (7) and with a given set of input data $\{f_{i}\}$. For sake of simplicity, in what follows we assume that the the form factor $f(z)$ is known for a series of distinct real values of $t=t_{i}<t_{+}$ (with i=1,2,,Ni=1,2,...,N), so that all the $z_{i}$ are real and different from each other. The untruncated BGL expansion (4) of the form factor $f(z)$ can be rewritten exactly in an additive form, as derived firstly in Refs. [38, 39, 40]. Here below, we use directly the final result of the computation described in Appendix B, namely
@@
The interpretation of the inequalities given by Eqs. (27-28) is as follows. The new inequality (27) is necessary to guarantee that the function (25) is consistent with unitarity and the remainder function (26) is analytic of the real type. Eq. (27) represents a filter acting directly on the input data and allows to select only the subset $\{\overline{f}_{i}\}$ of input data, which can be reproduced by BGL expansions satisfying the unitarity constraint (7). Furthermore, the inequality (28), acting on the coefficients $c_{k}$ of the remainder function $R(z)$, is a consequence of both the unitarity constraint $\chi[f]\leq\chi^{U}$, which is equivalent to the inequality (7) on the coefficients $a_{k}$ of the BGL expansion, and the new inequality (27). Indeed, an explicit calculation, shown in Appendix B, yields
@@
We stress that the above result implies two different unitarity constraints. The first one is given by $\chi[f]\leq\chi^{U}$, which implies the unitarity constraint (7), which guarantees that all the BGL expansions (24) are unitary. The other constraint is given by $\chi[\beta]\leq\chi^{U}$, which guarantees that the function (25) fulfills unitarity. It tells us whether a given set of input data $\{f_{i}\}$ is consistent with unitarity, i.e. whether it may be reproduced exactly by a BGL expansion (4) with coefficients $a_{k}$ satisfying the unitarity constraint (7). When Eq. (27) is not fulfilled, the input data $\{f_{i}\}$ are not consistent with unitarity and should be removed from any analysis since they contain non-unitary effects [^4] A consequence of the presence of non-unitary input data in a fitting procedure based on the BGL $z$-expansion is the value of the $\chi^{2}$-variable. Indeed, the latter one receives in general two contributions: the first one is related to the truncation order $M$ in Eq. (22), while the second one is due to the fact that non-unitary input data cannot be reproduced by a unitary BGL expansion. The latter contribution cannot be eliminated by increasing the order $M$ of the truncation and it is typically associated to a saturation of the dispersive bound $\chi^{U}$, i.e. $\chi[f]\simeq\chi^{U}$.. This may happen in various circumstances and we will describe some of them in a while. In all cases, starting from the given original sampling distribution for the input data, the application of the filter (27) allows to construct the subset of input data consistent with unitarity. In this respect, we mention the importance sampling procedure, developed in Ref. [36], which must be adopted in the case of a large number of input data points for a proper application of the unitarity filter $\chi[\beta]\leq\chi^{U}$.
We point out that, strictly speaking, the use of the unitarity filter (27) should not be viewed as a direct improvement of the BGL $z$-expansion (4) per se, but rather it represents an improvement of the standard unitary analysis based on the BGL $z$-expansion and applied to the reproduction of a given set of input data.
Thus, in order to summarize the results of this Section a fitting procedure based on a unitary BGL $z$-expansion of the form factor $f(z)$ should include the following steps.
@@
    Check the sample of input data $\{f_{i}\}$, generated according to a given original sampling distribution probability, against the unitarity filter $\chi[\beta]\leq\chi^{U}$.
@@
    Select only the subset of filtered input data $\{\overline{f}_{i}\}$ satisfying the unitarity constraint $\chi[\beta]\leq\chi^{U}$.
@@
    Apply the truncated BGL expansion (22) only to the unitary subset $\{\overline{f}_{i}\}$ and impose on the coefficients $a_{k}^{(M)}$ the unitarity constraint (23).
It often happens that a covariance matrix for the input data $\{f_{i}\}$ is provided or assumed for lacking of direct information. In these cases a sample of data can be generated according to a Gaussian multivariate distribution. We stress that our procedure of filtering the input data is not limited at all to the case of Gaussian statistics, but it can be applied equally well to any kind of initial sampling distribution probability. Analogously, the distribution of the filtered data $\{\overline{f}_{i}\}$ may contain deviations from a Gaussian behaviour, which can be properly taken into account by a direct use of the filtered subset of input data. In many phenomenological studies we have not observed sizable deviations from a normal statistics for the filtered subset $\{\overline{f}_{i}\}$ (see for instance Fig. 6 of Ref. [36]), while non-Gaussian effects may definitely occur for the (high-order) coefficients of the unitary BGL $z$-expansion. In any case it is recommended to present final results with error bars (and correlation matrix) only after checking that the distribution of the corresponding samples is basically Gaussian.
It is worthwhile to discuss briefly how non-unitary effects may plague the original input data $\{f_{i}\}$. We address explicitly few selected circumstances.
  • An interesting case is represented by the lattice calculations of the form factors relevant for the semileptonic $B\to D^{*}\ell\nu_{\ell}$ decays, performed in Refs. [7, 8, 9]. The synthetic data points provided by each of the various lattice collaborations are the results of simulations carried out for several gauge ensembles at fixed values of the lattice spacing and for heavy-quark masses below the physical $b$-quark mass. Then, the extrapolations to the continuum limit and to the physical $b$-quark mass were performed assuming a simple polynomial dependence of the (HQET-inspired) form factors upon the recoil variable. Therefore, there is no guarantee that the momentum dependences of the form factors with definite spin-parity satisfy the corresponding unitarity constraints. In Refs. [41, 14] we have applied to the lattice data of each collaboration the generalization of the unitarity filter (27) to the case of the $B\to D^{*}$ transition and we have found that only a limited subset of the input data satisfies unitarity.
  • Another important case is represented by a dataset coming from an experiment, like in the case of the electromagnetic form factor of the pion analyzed in Ref. [36]. Generally speaking, there is no reason why the systematic uncertainties should fulfill unitarity. Moreover, it may happen that statistical correlations among different data points are not addressed explicitly. The uncertainties provided for an experimental dataset may be viewed as a conservative estimate of the range of values of the form factor and, therefore, it should not be surprising if the application of the unitarity filtering selects a subset of the input data. Note that this does not necessarily imply the presence of effects due to operators beyond the Standard Model (SM). If the latter ones were instead present, part of them might lead to apparent violations of unitarity of the SM form factor. The investigation of such effects is well beyond the aim of the present paper and requires analyses at the level of cross sections or decay rates, where the $z$-expansion is not directly applicable.
  • In the case of multiple (experimental or theoretical) datasets, which are mutually uncorrelated, one can perform separate analyses for the individual input datasets and, after that, one can average in some way the individual results to compute a final prediction for the form factor (or for any observable of interest). Another option is to combine the different datasets in a unique larger dataset and check whether the final result for the form factor is consistent with the previous option. It is clear that independent fluctuations between different datasets may lead to violations of the unitarity constraint (23) for the coefficients of the BGL $z$-expansion. The way to properly avoid that is to apply our filtering procedure to the full dataset (see for instance Ref. [14]).
Since the application of the unitarity filter (27) can modify the sampling distribution probability for the input data, it is important to quantify and control its impact. To this end we introduce two quantities, $\Delta$ and $\epsilon$, defined as
$\Delta\equiv\left\{\frac{1}{N}\sum_{i,j=1}^{N}(\overline{f}_{i}-f_{i})C_{ij}^{-1}(\overline{f}_{j}-f_{j})\right\}^{1/2}~_{\overrightarrow{C_{ij}=\sigma_{i}^{2}\delta_{ij}}}~\left\{\frac{1}{N}\sum_{i=1}^{N}\frac{(\overline{f}_{i}-f_{i})^{2}}{\sigma_{i}^{2}}\right\}^{1/2}~,$~ (34)
and
$\epsilon\equiv\left\{\frac{1}{N}\sum_{i=1}^{N}\frac{\overline{\sigma}_{i}^{2}}{\sigma_{i}^{2}}\right\}^{1/2}~,$~ (35)
where $\{\overline{\sigma}_{i}\}$ represent the uncertainties of the “filtered” dataset, i.e. the ones obtained after the application of the unitarity filter(s) to the input dataset.
The quantity $\Delta$ represents the average deviation of the new values $\{\overline{f}_{i}\}$ from the starting ones $\{f_{i}\}$, measured with respect to the starting covariance matrix $C_{ij}$. In other words, $\Delta<1$ means that (on average) $\{\overline{f}_{i}\}$ deviates from $\{f_{i}\}$ by less than one standard deviation. The quantity $\epsilon$ is an estimate of how much the new uncertainties $\{\overline{\sigma}_{i}\}$ deviate (on average) from the original ones $\{\sigma_{i}\}$. It can be smaller or larger than unity depending on whether $\{\overline{\sigma}_{i}\}$ is (on average) smaller or larger than $\{\sigma_{i}\}$. If the application of the unitarity filter has no effect on the input dataset, one gets $\Delta=0$ and $\epsilon=1$.
Thus, in order to avoid an anomalous impact of the unitarity filter(s) on both the mean values and errors of the input dataset we may put an upper limit $\Delta^{*}$ to $\Delta$ and a lower limit $\epsilon^{*}$ to $\epsilon$, namely:
  • a)
    if $\Delta\leq\Delta^{*}$ and $\epsilon\geq\epsilon^{*}$, the“filtered” dataset can be kept and one can proceed with the computation of observables of interest for phenomenology;
  • b)
    if instead $\Delta>\Delta^{*}$ or $\epsilon<\epsilon^{*}$, one cannot proceed with a phenomenological study based on the filtered dataset.
Based on the values of $\Delta$ and $\epsilon$ found in the analyses of data on semileptonic decays of heavy mesons from Refs. [5, 42, 41, 43, 33, 14] and on the pion electromagnetic form factor [36], we argue that acceptable values of $\Delta^{*}$ and $\epsilon^{*}$ are $\Delta^{*}\simeq 1$ and $\epsilon^{*}\simeq 0.5$. In this way, if the condition $a)$ is satisfied, situations in which the average deviation of the new mean values of the FF exceed $1\sigma$ and the uncertainties are reduced by a factor of two or more are excluded. On the contrary, if the condition $b)$ is satisfied, one should avoid the use of the “filtered” dataset for phenomenological applications. We expect the results obtained implementing our filtering procedure to be consistent with those corresponding to other methodologies of imposing unitarity, like those already proposed and discussed in Refs. [38, 39, 40].
@@
which is nothing else than the final result of the DM method of Ref. [5], originally proposed in Refs. [27, 44]. Without relying on any truncation or fitting procedure, Eq. (37) represents the uniform distribution of values of the form factor $f(z)$ at a generic value of $z$ consistent with unitarity as well as with the input data $\{\overline{f}_{i}\}$. We stress that in the DM approach the unitarity filter $\chi[\beta]\leq\chi^{U}$ on the input data is always taken properly into account, since $\gamma(z)$ must be non-negative. The impact of the unitarity filter $\chi[\beta]\leq\chi^{U}$ has been addressed in the case of the semileptonic $B_{(s)}\to D_{(s)}^{(*)}\ell\nu_{\ell}$ decays in Refs.[5, 42, 41, 43, 14], for the semileptonic $B_{(s)}\to\pi(K)\ell\nu_{\ell}$ decays in Ref. [33] and for the pion electromagnetic form factor in Ref. [36]. We highlight that the inclusion of the unitarity filter $\chi[\beta]\leq\chi^{U}$ may lead to important reduction of the uncertainties associated to the extrapolation of the form factors, especially in presence of a large number of input data [^5] The changes in the mean values and errors on the FFs are always taken under control by evaluating the quantities $\Delta$ and $\epsilon$ given by Eqs. (34) and (35), respectively.. This is shown for instance in Fig. 6 of Ref. [14] in the case of the semileptonifc $B\to D^{*}\ell\nu_{\ell}$ form factors. There, the BGL results (green bands) are obtained without applying the unitary filtering (27) to the input data and, therefore, they differ from the DM predictions (the red bands) [^6] When the unitary BGL $z$-expansion is applied to the sample of unitarity-filtered input data, then the green bands coincide with the red ones, i.e. with the predictions of the DM approach.. We point out that an interesting generalization of Eq. (31), which includes also the first $K$ derivatives of the form factor at the origin $z=0$, was derived in Ref. [38] (see also Ref. [40]).
Finally, we comment briefly on the role of the auxiliary variable $t_{0}$, which fixes the value of $t$ at which $z(t_{0};t_{0})=0$. The value of $t_{0}$ may be chosen according to the requirement of accelerating as much as possible the convergence of the truncated BGL expansion (22). In Refs. [1, 3, 45] it is argued that for semileptonic process $H_{1}\to H_{2}\ell\nu_{\ell}$ an optimal choice could be $t_{0}=t_{opt}=(m_{1}+m_{2})\left(\sqrt{m_{1}}-\sqrt{m_{2}}\right)^{2}$, which minimizes the truncation error of the expansion (22) as estimated in Refs. [3, 4]. However, the form factor cannot depend on the choice of $t_{0}$. Since the Blaschke product $B(z)$ as well as the bound $\chi^{U}$ are independent of $t_{0}$ by construction, both the kinematical function $\phi(z)$ and the BGL coefficients $a_{k}$ must depend on $t_{0}$ in such a way to cancel any possible dependence of $f(z)$ on $t_{0}$. This implies that any difference in the truncated BGL $z$-expansions (22) corresponding to different values of $t_{0}$ may signal that the truncation error is not under full control.